OpenStudy (anonymous):

The function T(t) = 3t + 1, where 0

OpenStudy (anonymous):

integral is 33/2 divide by 3 get 11/2

OpenStudy (anonymous):

wow before i was so rudely interrupted. \[\frac{1}{3}\int_0^3 3t+1 dt\]

OpenStudy (anonymous):

what do you mean satellite?

OpenStudy (anonymous):

you want average value yes?

OpenStudy (anonymous):

average value of a function over an interval [a,b] is \[\frac{1}{b-a}\int_a^b f(t)dt\]

OpenStudy (anonymous):

in your case noon is 0, three o'clock is 3. average is \[\frac{1}{3}\int_0^3 (3t+1)dt\]

OpenStudy (anonymous):

the integral is easy to solve. it is \[\frac{33}{2}\]

OpenStudy (anonymous):

divide by 3 give \[\frac{11}{2}\]so average is 5.2 must have been a cold day

OpenStudy (anonymous):

sorry i mean 5.5

OpenStudy (anonymous):

now the second part of the question asks when the temperature was in fact 5.5 so solve for t [\3t+1=5.5\] \[3t=4.5\] \[t=1.5\]not coincidentally half way between 0 and 3

OpenStudy (anonymous):

half way because you have a line

OpenStudy (anonymous):

The animosity is so strong amongst these comments. Thanks.

OpenStudy (anonymous):

didn't mean any sorry

OpenStudy (anonymous):

the 'before i was rudely interrupted" part is because the site crashed right as i was typing. it was out for an hour or so

OpenStudy (anonymous):

Oh. It was just a misunderstanding (: Thanks again